# Fair Foundation Models for Medical Image Analysis: Challenges and Perspectives

### v1

@@ Abstract
**footnotetext: Correspondence: dilermando.queiroz@unifesp.br
@@ 1 Introduction
Figure 1: Conceptual framework. Sequential phases of FMs development that structure this work.
AI in medicine offers transformative potential through improving access to diagnostics and enhancing the speed and quality of medical care 1, 2. AI tools extend healthcare services, particularly in resource-limited regions, thereby making care more efficient and accessible 3. However, these advancements raise critical ethical concerns that emphasize the need for a fair and equitable distribution of benefits across all populations 4, 5, 6. We must develop and apply these technologies responsibly to uphold bioethical principles of justice, autonomy, beneficence, and non-maleficence to prevent discrimination and promote inclusive healthcare for all 7.
Governments worldwide are establishing regulatory frameworks for artificial intelligence across critical sectors. The EU AI Act 8, the first comprehensive regulation of its kind, has introduced a risk-based classification system for AI applications. In parallel, the U.S. Office for Civil Rights has enacted specific protections against algorithmic bias in healthcare through the Affordable Care Act 9. These regulatory initiatives reflect a coordinated global effort to implement evidence-based guidelines that ensure the fairness, safety, and equity of AI systems 10, 11, 12, 13, 14, 15, 16, 17.
FMs serve as essential components in AI by enabling diverse tasks across text, image, video, audio, and other domains through their versatility and scalability 18. Through training on massive datasets, these models capture broad patterns within and across domains. However, significant challenges persist in developing ethical models that can effectively identify and reduce inherent biases 19, 20, 21. While bias persists across AI systems, FMs demonstrate significant potential for their mitigation 22, 23, thus creating opportunities for unified models that equitably serve diverse populations while driving greater inclusion in AI.
The production of FMs requires substantial resources, including specialized labor, large datasets, and significant computational infrastructure. These high costs restrict development capabilities to select countries, institutions, and companies, thereby increasing the risk of global inequalities. Stakeholders in regions such as Africa have indicated their inability to develop such models, which potentially widens the disparity between populations that benefit from AI and those that do not 24. Addressing this challenge requires developing strategies that facilitate the creation of globally accessible FMs while enabling broad participation in their production.
@@
A substantial body of research has empirically demonstrated inherent biases and established initial frameworks for evaluating fairness in these systems 25, 20, 19, 21. While comprehensive investigations of trustworthiness in medical imaging FMs have been conducted 26, these studies primarily address fairness as one component within broader trustworthiness considerations. Additionally, existing analyses provide thorough technical perspectives on medical imaging FMs 2, 27 but lack systematic examination of fairness considerations. Current reviews of AI fairness in healthcare 28, 29, 30, 5, 31 have not adequately addressed the distinct challenges posed by FMs. To our knowledge, no study has comprehensively reviewed the field regarding challenges and approaches to ensuring fairness in FMs, underscoring the need for systematic interventions throughout their development lifecycle.
@@ 2 Background and Taxonomy
This section delineates the fundamental principles underlying fairness and foundation models.
@@ 2.1 Fairness
Principles of Trustworthy AI. The development of medical image analysis AI systems, to achieve trustworthiness, is guided by six fundamental principles: fairness, universality, traceability, usability, robustness, and explainability 13. Among these, fairness, which ensures non-discriminatory outcomes across diverse patient populations, represents a critical determinant of ethical AI implementation. Our analysis centers on fairness and examines its interdependencies with other principles, specifically how it interacts with robustness to ensure reliable system performance and with traceability to maintain systematic accountability, thus demonstrating how these principles collectively contribute to trustworthy AI in clinical settings.
Healthcare Disparities. Differences in national healthcare delivery capacities result in variable health outcomes across populations 32, 33. Within countries that maintain public healthcare systems, disparities persist because of multiple factors, including race, gender, age, ethnicity, body mass index, education, insurance status, and geographic location 34, 35. These determinants influence an individual’s capacity to access treatment, obtain quality care, and achieve favorable health outcomes.
AI Bias Manifestation. AI models frequently assimilate and reproduce biases inherent in their training data, subsequently reflecting and intensifying societal inequalities. These models often depend on spurious correlations, employing computational shortcuts that amplify existing biases 36, 37, 38. A well-designed fair algorithm produces impartial decisions by ensuring equitable outcomes across demographic groups without discrimination based on sensitive attributes such as race, gender, or age. Despite the implementation of controlled datasets and balanced groups, bias can manifest through various factors, including image complexity and labeling inconsistencies 39. Consequently, unintended biases may emerge even under optimal conditions, highlighting the necessity for continuous evaluation and mitigation procedures to maintain algorithmic fairness.
Bias Mitigation Strategies. Previous research has established three primary classifications for bias mitigation strategies: pre-processing, in-processing, and post-processing 29, 28, 30, 31. Pre-processing approaches modify datasets through demographic representation balancing, sensitive feature removal (such as race and gender), or synthetic data augmentation to enhance diversity 40, 41, 42. In-processing methods alter the training process by integrating fairness constraints into the loss function or implementing adversarial training to prevent bias acquisition 43, 44, 45. Post-processing techniques adjust model predictions after training through methods such as prediction calibration to satisfy fairness criteria 46, 47, 48.
@@ 2.2 Foundation Models
Technical Foundations. FMs are trained on extensive datasets that can be efficiently adapted to multiple downstream tasks through fine-tuning, thereby eliminating the necessity of training specialized models from scratch. The Vision Transformer (ViT) 49 serves as a leading FMs architecture in computer vision for extracting fundamental image features. The integration of self-supervised learning techniques, such as Masked Autoencoder (MAE) 50 and contrastive learning methods like SimCLR 51, enables these models to learn directly from large volumes of unlabeled data, thus largely mitigating the need for costly manual labeling processes. The Image-based Joint-Embedding Predictive Architecture (I-JEPA) introduces a self-supervised learning approach that predicts semantic embeddings directly, offering enhanced efficiency compared to pixel-reconstruction methods 52.
@@
Capabilities and Applications. Through the utilization of massive unlabeled datasets, advanced architectures, and self-supervised learning techniques, FMs acquire comprehensive image representations within their training domains. The fine-tuning process facilitates task-specific adaptation through minimal adjustments and limited labeled data requirements. Several key capabilities establish their essential role in scalable, real-world applications: rapid task adaptation, robustness to distribution shifts, and efficient utilization of labeled data 53, 54, 55, 56.
Multimodal Integration. These models facilitate training across diverse data types within a unified architectural framework. Vision-Language Models (VLMs) 57 demonstrate this capability by integrating visual and linguistic processing within a single computational structure. This architectural integration enables simultaneous interpretation and generation of information across multiple modalities, thereby establishing new possibilities for applications that require image and textual data.
@@
Domain Specialization. Domain-specific FMs constitute specialized architectures tailored to address the distinctive characteristics of particular domains, ranging from natural image processing 58 to medical imaging applications 55, 54. In contrast to general-purpose models, these specialized frameworks concentrate on the unique data structures and analytical challenges inherent to their respective domains. Medical imaging presents a notable example, where image characteristics vary substantially in resolution, ranging from whole-organ visualization to cellular-level structures, with data originating from diverse acquisition modalities including X-ray radiography, computed tomography (CT), magnetic resonance imaging (MRI), and ultrasonography.
Medical FMs Categories. The inherent diversity of medical imaging modalities presents significant challenges in developing a comprehensive global model. To address this complexity, researchers have established a systematic categorization of medical image FMs into distinct segments: general Medical Image FMs for broad applications, Modality-specific FMs that address particular imaging techniques, Organ-specific FMs that target anatomical structures, and Task-specific FMs that focus on specialized medical procedures such as segmentation or classification 2.
@@ 3 Data Documentation
Figure 2: Global distribution of medical imaging data: Geographic visualization of dataset volumes across countries (excluding datasets from Table 1 that span multiple locations).
@@
Dataset Model Region Modality Images Demographics Origin License
Moorfields BioResource 001 59 Vision Ocular Multimodal 26,548,820 A UK Restricted
MedTrinity-25M60 Multimodal Multiple Multimodal 25,000,000 Multiple Group
PMC-15M61 Multimodal Multiple Multimodal 15,000,000 Multiple CC BY-SA
BRATS24-MICCAI62 Vision Cerebral MRI 2,535,132 USA CC BY 4.0
PMC-OA63 Multimodal Multiple Multimodal 1,600,000 A Multiple OpenRAIL
RadImageNet64 Vision Multiple Multimodal 1,350,000 USA CC BY 4.0
ISIC65 Vision Dermal Dermoscopy 1,162,456 A, S Multiple CC BY-NC-SA
TCGA66 Vision Dermal Histology 1,142,221 A, R, S USA CC BY-NC-SA
HyperKvasir67 Vision Colon Endoscopy 1,000,000 Norway CC BY 4.0
BRATS-ISBI68 Vision Cerebral MRI 987,340 Multiple CC BY 4.0
BHX69 Vision Cerebral MRI 973,908 India CC BY 4.0
LDPolypVideo70 Vision Colon Endoscopy 901,666 China
MIMIC-CXR-JPG71 Multimodal Pulmonary Radiograph 370,955 A, R, S USA PHDL
CheXpert72 Multimodal Pulmonary Radiograph 222,793 A, R, S USA RUA
PadChest73 Multimodal Pulmonary Radiograph 160868 A, S Spain RUA
SUN-SEG74 Vision Colon Endoscopy 158,690 A, S Japan MIT
NIH-CXR1475 Vision Pulmonary Radiograph 112,120 A, S USA CC0
Kermany et al.76 Vision Ocular OCT 108,312 USA CC BY 4.0
EyePACS77 Vision Ocular Fundus 92,501 India CC BY 4.0
BRAX78 Vision Pulmonary Radiograph 40,967 A, S Brazil PHDL
MURA79 Vision Multiple Radiograph 40,561 USA CC BY 4.0
ASU-Mayo80 Vision Colon Endoscopy 19,400 USA
BRSET81 Vision Ocular Fundus 16,266 A, S, N Brazil PHDL
Harvard-FairVLMed82 Multimodal Ocular SLO 10,000 A, R, S, E USA CC BY-NC-SA
@@ 3.1 Data Creation
The first step in FMs development involves comprehensive data collection. Healthcare facilities generate medical images through specialized imaging equipment that captures various anatomical structures. The pre-training phase utilizes extensive unlabeled datasets, enabling FMs to acquire general representations and discover underlying patterns without annotation requirements. For subsequent fine-tuning processes, labeled data plays an essential role in enabling models to develop task-specific capabilities through the refinement of learned representations. Both unlabeled and labeled datasets must maintain high quality and diversity standards, as these characteristics fundamentally influence FMs performance and generalization capabilities 83, 58.
@@
Criteria for Inclusion. Our investigation of large-scale datasets employed in FMs followed a systematic approach, establishing a minimum threshold of 10,000 images for inclusion (Table 1). The geographic distribution of dataset origins appears in Figure 2. While not exhaustive, this methodological approach enabled the systematic curation of 74,184,415 medical images, representing a comprehensive cross-section of diverse imaging modalities and anatomical regions. The comprehensive review encompassed unimodal imaging and multimodal image-text paired datasets, spanning diverse imaging modalities and anatomical regions. This methodological documentation framework facilitates a systematic assessment of the current medical imaging data ecosystem while elucidating opportunities for enhanced demographic representation in FMs development.
@@
Pre-training. The utilization of large unlabeled datasets for FMs training significantly reduces dependence on costly and time-intensive labeled data collection. Advanced learning methodologies, particularly self-supervised learning approaches 51, 50, 52, facilitate effective utilization of unlabeled data while minimizing labeling bias. This methodology proves especially valuable in medical contexts, where individual clinician annotation preferences, influenced by patient attributes, may introduce systematic biases 84. Using unlabeled data enables FMs to identify generalizable patterns while simultaneously reducing both cost constraints and annotation-induced biases.
@@
The development of robust datasets for FMs presents significant challenges in ensuring comprehensive representation across populations, imaging modalities, equipment specifications, and disease classifications. The accumulation of large-scale data frequently amplifies existing biases and imbalances that reflect underlying disparities in healthcare access and infrastructure. As illustrated in Figure 2, the predominant source of available datasets resides in developed nations, which substantially constrains the diversity of documented diseases and patient demographics. This geographic concentration of data resources is particularly evident in regions such as Africa, where large-scale medical imaging datasets remain notably absent.
@@
Fine-tuning. Smaller, more accurate, and more representative datasets are ideal for specializing FMs for specific tasks. This specialization is achieved through supervised learning on labeled datasets. Due to their data efficiency and generalization capabilities 53, 54, 55, 56, FMs require less data than training supervised models from scratch. Consequently, fine-tuning FMs offers significant advantages, as it requires fewer computational resources and enables countries and institutions to adapt these models to their specific tasks with greater accessibility.
@@
Precise labels and comprehensive patient information, including gender, age, and race, facilitate the evaluation of model performance and enable systematic bias identification 78, 81, 72, 85. This demographic information proves essential for model evaluation, as it enables more rigorous identification and mitigation of potential biases (discussed in the evaluation section 4.2). The availability of labeled data supports dataset balancing across specific attributes, thereby enhancing model fairness and performance. This represents a significant advantage over unlabeled data, which presents greater challenges in achieving these objectives 86.
@@
Generative Models. Generative models constitute a driving force behind recent advancements in AI and its applications. These models learn to approximate the probability distribution of output features conditioned on input features, enabling the generation of novel data instances that closely approximate the characteristics of the training data. Advanced techniques, including generative adversarial networks (GANs) 87, variational autoencoders (VAEs) 88, and diffusion models 89, have significantly enhanced the capabilities of generative AI systems.
@@
Data augmentation through generative approaches facilitates both pre-training and fine-tuning stages. These models enable dataset balancing across sensitive attributes through synthetic data generation for underrepresented patient groups while simultaneously enhancing model robustness through the generation of challenging cases 39. Empirical studies demonstrate that fine-tuning models with augmented data improve both fairness and robustness, resulting in enhanced performance across diverse populations 90. However, generative models may inadvertently amplify biases present in training data, such as generating synthetic medical images that systematically underrepresent darker skin tones 91, which necessitates robust bias mitigation strategies when utilizing synthetic data for augmentation.

### v2

@@ Abstract
Keywords Foundation Models $\cdot$ Fairness $\cdot$ World Health
@@ 1 Introduction
Figure 1: Conceptual framework. This figure delineates sequential phases in FMs development, illustrating principal challenges for achieving fairness. Three focal domains are emphasized: data documentation, curation dataset ensuring diversity and quality to detect and mitigate bias; environmental impact, spanning training, deployment, and resource efficiency; and policymakers, establish governance, standards, and resource distribution to ensure ethical, equitable FMs access.
AI in medicine offers transformative potential through improving access to diagnostics and enhancing the speed and quality of medical care [110, 169]. AI tools extend healthcare services, particularly in resource-limited regions, thereby making care more efficient and accessible [151]. However, these advancements raise critical ethical concerns that emphasize the need for a fair and equitable distribution of benefits across all populations [125, 129, 104]. We must develop and apply these technologies responsibly to uphold bioethical principles of justice, autonomy, beneficence, and non-maleficence to prevent discrimination and promote inclusive healthcare for all [13].
Governments worldwide are establishing regulatory frameworks for AI across critical sectors. The EU AI Act [46], the first comprehensive regulation of its kind, has introduced a risk-based classification system for AI applications. In parallel, the U.S. Office for Civil Rights has enacted specific protections against algorithmic bias in healthcare through the Affordable Care Act [143]. These regulatory initiatives reflect a coordinated global effort to implement evidence-based guidelines that ensure the fairness, safety, and equity of AI systems [63, 94, 116, 85, 145, 34, 139, 92].
FMs serve as essential components in AI by enabling diverse tasks across text, image, video, audio, and other domains through their versatility and scalability [16]. Through training on massive datasets, these models capture broad patterns within and across domains. However, significant challenges persist in developing ethical models that can effectively identify and reduce inherent biases [77, 53, 88]. While bias persists across AI systems, FMs demonstrate significant potential for their mitigation [144, 123], thus creating opportunities for unified models that equitably serve diverse populations while driving greater inclusion in AI.
The production of FMs requires substantial resources, including specialized labor, large datasets, and significant computational infrastructure. These high costs restrict development capabilities to select countries, institutions, and companies, thereby increasing the risk of global inequalities. Stakeholders in regions such as Africa have indicated their inability to develop such models, which potentially widens the disparity between populations that benefit from AI and those that do not [2]. Addressing this challenge requires developing strategies that facilitate the creation of globally accessible FMs while enabling broad participation in their production.
@@
A substantial body of research has empirically demonstrated inherent biases in FMs and established initial frameworks for fairness evaluation [71, 53, 77, 88]. The FairMedFM benchmark, for instance, provides a standardized basis for these empirical assessments [71]. While comprehensive investigations of trustworthiness in medical imaging FMs have been conducted [133], these studies primarily address fairness as one component within broader trustworthiness considerations. Additionally, existing analyses provide thorough technical perspectives on medical imaging FMs [169, 87] but lack systematic examination of fairness considerations. Current reviews of AI fairness in healthcare [30, 44, 105, 129, 161] have not adequately addressed the distinct challenges posed by FMs. To our knowledge, no study has adopted a broader conceptual lens to address the interconnected ethical, environmental, and governance challenges in ensuring fairness in FMs. To address this gap, this review investigates the development lifecycle from a narrative perspective, identifying key opportunities and challenges for fair FMs.
@@ 2 Background and Taxonomy
This section delineates the fundamental principles underlying fairness and FMs. We also present a table summarizing key concepts and strategies relevant to our review (Table 1).
Table 1: Taxonomy. Key concepts and strategies encompassing fairness, protected attributes, foundation models, vision-language models, and learning approaches relevant to scalable and equitable AI systems.
Term Summary Key Concepts
Fairness The principle that ensures equitable treatment and non-discriminatory outcomes across diverse groups defined by protected attributes in algorithmic decision-making. Equity, bias mitigation, performance parity, protected groups.
Hallucination The generation of plausible but false or fabricated information not grounded in real-world facts. Misinformation, clinical risk, reliability, uncertainty, inaccurate knowledge.
Protected Attributes Patient demographic and clinical metadata that require safeguarding to prevent bias in algorithmic decision-making, ensuring fairness across groups defined by these attributes (e.g., age, gender). Protected attributes, bias prevention, demographic metadata, group fairness.
Foundation Models (FMs) Large-scale pretrained models adaptable to diverse tasks, with emphasis on self-supervised and weakly-supervised training strategies. Pre-training, downstream task, scalability, robustness.
Vision Language Models (VLMs) Models that integrate visual and linguistic data processing, typically trained using weakly supervised learning to leverage large-scale paired image-text data. Multimodal learning, weakly supervised learning, visual-linguistic integration, FMs.
World Models Models that simulate and predict the dynamics of the real world, enabling anticipation of future states and informed decision-making. Environment Simulation, latent space prediction, planning.
Generative Models Models that learn data patterns to create new, original content resembling the training data. Content generation, hallucinations, synthetic data.
Knowledge Agglomeration Knowledge agglomeration synthesizes knowledge by distilling diverse representations from multiple FMs into a single, unified agglomerative model. Knowledge distillation, Model merging, agglomerative models.
Self-supervised Learning Training method using large unlabeled data by leveraging inherent data structures. Unlabeled data, Masked Autoencoder (MAE) [60], Simple Framework for Contrastive Learning of Visual Representations (SimCLR) [32], Image-based Joint-Embedding Predictive Architecture (I-JEPA) [8].
Weakly-supervised Learning Learning approach that leverages limited, noisy, or imprecise labels to guide model training, commonly used in large-scale and multimodal datasets. Noisy labels, limited supervision, semi-supervised methods, VLMs, Contrastive Language-Image Pretraining (CLIP) [124], Sigmoid Loss for Language Image Pre-Training (SigLIP) [165].
@@ 2.1 Fairness
Principles of Trustworthy AI. The development of medical image analysis AI systems, to achieve trustworthiness, is guided by six fundamental principles: fairness, universality, traceability, usability, robustness, and explainability [85]. Among these, fairness, which ensures non-discriminatory outcomes across diverse patient populations, represents a critical determinant of ethical AI implementation. Our analysis centers on fairness and examines its interdependencies with other principles, specifically how it interacts with robustness to ensure reliable system performance and with traceability to maintain systematic accountability, thus demonstrating how these principles collectively contribute to trustworthy AI in clinical settings. We specifically center hallucinations as a critical reliability issue, as these generated false but plausible outputs can undermine fairness by causing biased or unequal clinical outcomes.
Healthcare Disparities. Differences in national healthcare delivery capacities result in variable health outcomes across populations [154, 28]. Within countries that maintain public healthcare systems, disparities persist because of multiple factors, including race, gender, age, ethnicity, body mass index, education, insurance status, and geographic location [11, 152]. These determinants influence an individual’s capacity to access treatment, obtain quality care, and achieve favorable health outcomes.
AI Bias Manifestation. AI models frequently assimilate and reproduce biases inherent in their training data, subsequently reflecting and intensifying societal inequalities. These models often depend on spurious correlations, employing computational shortcuts that amplify existing biases [50, 177, 52]. A well-designed fair algorithm produces impartial decisions by ensuring equitable outcomes across demographic groups without discrimination based on protected attributes such as race, gender, or age. Despite the implementation of controlled datasets and balanced groups, bias can manifest through various factors, including image complexity and labeling inconsistencies [36]. Consequently, unintended biases may emerge even under optimal conditions, highlighting the necessity for continuous evaluation and mitigation procedures to maintain algorithmic fairness.
Bias Mitigation Strategies. Previous research has established three primary classifications for bias mitigation strategies: pre-processing, in-processing, and post-processing [44, 30, 105, 161]. Pre-processing approaches modify datasets through demographic representation balancing, protected feature removal (such as race and gender), or synthetic data augmentation to enhance diversity [21, 115, 23]. In-processing methods alter the training process by integrating fairness constraints into the loss function or implementing adversarial training to prevent bias acquisition [25, 153, 163]. Post-processing techniques adjust model predictions after training through methods such as prediction calibration to satisfy fairness criteria [73, 122, 33].
@@ 2.2 Foundation Models
FMs are trained on extensive datasets that can be efficiently adapted to multiple downstream tasks through fine-tuning, thereby eliminating the necessity of training specialized models from scratch. The Vision Transformer (ViT) [43] serves as a leading FMs architecture in computer vision for extracting fundamental image features. The integration of self-supervised learning techniques, such as MAE [60] and contrastive learning methods like SimCLR [32], enables these models to learn directly from large volumes of unlabeled data, thus largely mitigating the need for costly manual labeling processes. The I-JEPA introduces a self-supervised learning approach that predicts semantic embeddings directly, offering enhanced efficiency compared to pixel-reconstruction methods [8].
@@
Capabilities and Applications. Through the utilization of massive unlabeled datasets, advanced architectures, and self-supervised learning techniques, FMs acquire comprehensive image representations within their training domains. The fine-tuning process facilitates task-specific adaptation through minimal adjustments and limited labeled data requirements. Several key capabilities establish their essential role in scalable, real-world applications: rapid task adaptation, robustness to distribution shifts, and efficient utilization of labeled data [174, 9, 155, 142]. For example, FMs improve diagnostic accuracy by 11.5% relative to supervised methods and, in out-of-distribution settings, achieve comparable performance using only 1–33% of the labeled data [9].
Multimodal Integration. These models facilitate training across diverse data types within a unified architectural framework. VLMs demonstrate this capability by integrating visual and linguistic processing through weakly supervised learning that combines image captions with contrastive losses, such as CLIP [124]. This integration supports simultaneous interpretation and generation of information across modalities. Subsequent work has improved training by replacing contrastive loss with a sigmoid-based loss, as in SigLIP [166], while other studies highlight the crucial role of data quality over model architecture or pretraining objectives [160].
@@
Domain Specialization. Domain-specific FMs constitute specialized architectures tailored to address the distinctive characteristics of particular domains, ranging from natural image processing [117], medical imaging applications [155, 9] to VLMs tailored for medical domain [149, 132]. In contrast to general-purpose models, these specialized frameworks concentrate on the unique data structures and analytical challenges inherent to their respective domains. Medical imaging presents a notable example, where image characteristics vary substantially in resolution, ranging from whole-organ visualization to cellular-level structures, with data originating from diverse acquisition modalities including X-ray radiography, computed tomography (CT), magnetic resonance imaging (MRI), and ultrasonography [169].
@@ 3 Data Documentation
Figure 2: Global distribution of medical imaging data: This geographic visualization depicts the volume of medical imaging datasets by country, excluding multi-country datasets listed in Table 2. The figure highlights pronounced disparities in data representation, underscoring its critical role in the development of equitable FMs. Notably, a limited number of countries account for the majority of available medical imaging data.
@@
Dataset Model Language Region Modality Images Demographics Origin License
Moorfields BioResource 001 [49] Vision Ocular Multimodal 26,548,820 A UK Restricted
MedTrinity-25M[158] Multimodal English Multiple Multimodal 25,000,000 Multiple Group
PMC-15M[170] Multimodal English Multiple Multimodal 15,000,000 Multiple CC BY-SA
BRATS24-MICCAI[37] Vision Cerebral MRI 2,535,132 USA CC BY 4.0
PMC-OA[89] Multimodal English Multiple Multimodal 1,600,000 A Multiple OpenRAIL
RadImageNet[106] Vision Multiple Multimodal 1,350,000 USA CC BY 4.0
ISIC[164] Vision Dermal Dermoscopy 1,162,456 A, S Multiple CC BY-NC-SA
TCGA[75] Vision Dermal Histology 1,142,221 A, R, S USA CC BY-NC-SA
HyperKvasir[20] Vision Colon Endoscopy 1,000,000 Norway CC BY 4.0
BRATS-ISBI[74] Vision Cerebral MRI 987,340 Multiple CC BY 4.0
BHX[128] Vision Cerebral MRI 973,908 India CC BY 4.0
LDPolypVideo[100] Vision Colon Endoscopy 901,666 China
MIMIC-CXR-JPG[72] Multimodal English Pulmonary Radiograph 370,955 A, R, S USA PHDL
CheXpert[67] Multimodal English Pulmonary Radiograph 222,793 A, R, S USA RUA
PadChest[22] Multimodal Spanish Pulmonary Radiograph 160,868 A, S Spain RUA
SUN-SEG[68] Vision Colon Endoscopy 158,690 A, S Japan MIT
NIH-CXR14[147] Vision Pulmonary Radiograph 112,120 A, S USA CC0
Kermany et al.[76] Vision Ocular OCT 108,312 USA CC BY 4.0
EyePACS[58] Vision Ocular Fundus 92,501 India CC BY 4.0
BRAX[127] Vision Pulmonary Radiograph 40,967 A, S Brazil PHDL
MURA[126] Vision Multiple Radiograph 40,561 USA CC BY 4.0
ASU-Mayo[137] Vision Colon Endoscopy 19,400 USA
BRSET[113] Vision Ocular Fundus 16,266 A, S, N Brazil PHDL
Harvard-FairVLMed[97] Multimodal English Ocular SLO 10,000 A, R, S, E USA CC BY-NC-SA
@@ 3.1 Data Creation
The first step in FMs development involves comprehensive data collection. Healthcare facilities generate medical images through specialized imaging equipment that captures various anatomical structures. The pre-training phase utilizes extensive unlabeled datasets, enabling FMs to acquire general representations and discover underlying patterns without annotation requirements. For subsequent fine-tuning processes, labeled data plays an essential role in enabling models to develop task-specific capabilities through the refinement of learned representations. Both unlabeled and labeled datasets must maintain high quality and diversity standards, as these characteristics fundamentally influence FMs performance and generalization capabilities [39, 117].
Inclusion Criteria. Datasets were identified for inclusion in Table 2 through a structured, multi-step process. The initial step involved identifying the pre-training datasets for the principal FMs listed in a recent survey [133]. To enhance geographic representation, a supplementary search was conducted on the Hugging Face and PhysioNet platforms to identify datasets from underrepresented regions. A minimum size threshold of 10,000 images was applied to all datasets to ensure they represent large-scale collections. Finally, metadata for each selected dataset were manually extracted from the corresponding publication and official website to populate the table.
@@
While not exhaustive, this methodological approach enabled the systematic curation of 74,184,415 medical images, representing a comprehensive cross-section of diverse imaging modalities and anatomical regions. The comprehensive review encompassed unimodal imaging and multimodal image-text paired datasets, spanning diverse imaging modalities and anatomical regions. This methodological documentation framework facilitates a systematic assessment of the current medical imaging data ecosystem while elucidating opportunities for enhanced demographic representation in FMs development.The geographic distribution of dataset origins appears in Figure 2.
@@
Pre-training. The utilization of large unlabeled datasets for FMs training significantly reduces dependence on costly and time-intensive labeled data collection. Advanced learning methodologies, particularly self-supervised learning approaches [32, 60, 8], facilitate effective utilization of unlabeled data while minimizing labeling bias. This methodology proves especially valuable in medical contexts, where individual clinician annotation preferences, influenced by patient attributes, may introduce systematic biases [15]. Using unlabeled data enables FMs to identify generalizable patterns while simultaneously reducing both cost constraints and annotation-induced biases.
@@
The development of robust datasets for FMs presents significant challenges in ensuring comprehensive representation across populations, imaging modalities, equipment specifications, and disease classifications. The accumulation of large-scale data frequently amplifies existing biases and imbalances that reflect underlying disparities in healthcare access and infrastructure. As illustrated in Figure 2, the predominant source of available datasets resides in developed nations, which substantially constrains the diversity of documented diseases and patient demographics. This geographic concentration of data resources is particularly evident in regions such as Africa, where large-scale medical imaging datasets remain notably absent (see Section 4.1 for a detailed discussion of the implications of such underrepresentation).
@@
Fine-tuning. Smaller, more accurate, and more representative datasets are ideal for specializing FMs for specific tasks. This specialization is achieved through supervised learning on labeled datasets. Due to their data efficiency and generalization capabilities [174, 9, 155, 142], FMs require less data than training supervised models from scratch. Consequently, fine-tuning FMs offers significant advantages, as it requires fewer computational resources and enables countries and institutions to adapt these models to their specific tasks with greater accessibility.
@@
Precise labels and comprehensive patient information, including gender, age, and race, facilitate the evaluation of model performance and enable systematic bias identification [127, 113, 67, 56]. This demographic information proves essential for model evaluation, as it enables more rigorous identification and mitigation of potential biases (discussed in the evaluation section 4.2). The availability of labeled data supports dataset balancing across specific attributes, thereby enhancing model fairness and performance. This represents a significant advantage over unlabeled data, which presents greater challenges in achieving these objectives [7].
@@
Generative Models. Generative models constitute a driving force behind recent advancements in AI and its applications. These models learn to approximate the probability distribution of output features conditioned on input features, enabling the generation of novel data instances that closely approximate the characteristics of the training data. Advanced techniques, including generative adversarial networks (GANs) [93], variational autoencoders (VAEs) [120], diffusion models [121], have significantly enhanced the capabilities of generative AI systems.
@@
Data augmentation through generative approaches facilitates both pre-training and fine-tuning stages. These models enable dataset balancing across protected attributes through synthetic data generation for underrepresented patient groups and enhance model robustness by generating challenging cases, reducing the fairness gap in chest radiograph classifiers trained on synthetic and real images by 44.6% [36]. Empirical studies demonstrate that fine-tuning models with augmented data improve both fairness and robustness, resulting in enhanced performance across diverse populations [81]. However, generative models may inadvertently amplify biases present in training data, such as generating synthetic medical images that systematically underrepresent darker skin tones [168], which necessitates robust bias mitigation strategies when utilizing synthetic data for augmentation.
Hallucinations. Understood as model-generated content that is not aligned with the real world, constitute a major source of misinformation that undermines the reliability of text and image generation models. They are typically associated with high uncertainty and incomplete or inaccurate knowledge [91, 69, 65, 101], particularly in instances where the model lacks relevant training data and exhibits low confidence [79, 157, 91]. As illustrated in Figure 2, large regions of the world are underrepresented in current datasets, which implies that models are more prone to hallucinate in these contexts because their underlying knowledge is partial or missing [31]. Nevertheless, it is important to note that hallucinations also occur in settings where the model has the knowledge and model confidence is high [134], indicating that they remain a persistent concern even under apparently well-specified conditions.